OpenStudy (kewlgeek555):

MEDAL + TESTIMONIAL Part B: The function f(h) shown below represents the number of tennis matches Diana won each month when she practiced tennis for h hours every day. If f(h) = 3h + 1 Calculate and interpret the meaning of f(3). (4 points) Part C: Write an ordered pair to represent the input and output of the function in Part B when Diana practices tennis for b hours every day. (2 points)

OpenStudy (kewlgeek555):

I have Part B figured out. I just don't know Part C. If you need my work for Part B, tell me.

OpenStudy (ranga):

Tell me what you got for Part B

OpenStudy (kewlgeek555):

Okay. This is what I put: Part B: Well, if you change the h to 3, you have to change all h to 3, so the new equation would loo like this: f(3) = 3 * 3 + 1 Now, usually, because the whole f(3) thing confuses me, I remind myself that f(3) = h. So, I just solve if I were solving for h. y = 3 * 3 + 1 y = 9 + 1 y = 10 f(3) = 10

OpenStudy (ranga):

Part B is correct. You found f(3) correctly. But they also wanted you to interpret the meaning of f(3) = 10.

OpenStudy (kewlgeek555):

Hm...I don't know about that. I think that he played tennis for 10 hours?

OpenStudy (ranga):

f(h) = 3h + 1 For part C, evaluate f(b) = ? just put h = b in f(h) = 3h + 1

OpenStudy (kewlgeek555):

Really, @LanaPhillips14 it seems that you do not know how to leave someone alone.

OpenStudy (kewlgeek555):

So, I just re-write the equation?

OpenStudy (anonymous):

omg

OpenStudy (kewlgeek555):

Please stop distracting Lana.

OpenStudy (ranga):

In f(h) = 3h + 1 h is hours of tennis Diana practiced each day and f(h) is the number of matches Diana won that month. So h = 3, f(3) = 10 means id Diana practiced 3 hours of tennis each day and she won 10 matches that month.

OpenStudy (kewlgeek555):

Okay, now I get it - thank you! I have one more.

OpenStudy (kewlgeek555):

A truck weighs 4500 kg. The truck is packed with bales of wool, each weighing 150 kg. Part A: Write a function to show the total weight of the truck f(w) when it is loaded with n bales of wool. (4 points)

OpenStudy (ranga):

For Part C, put h = b in f(h) = 3h + 1 f(b) = 3b + 1 Put the answer as an ordered pair. So the answer is (b, 3b+1)

OpenStudy (kewlgeek555):

Okay. Can you answer my other question?

OpenStudy (ranga):

Does the second sentence say f(w) ?

OpenStudy (kewlgeek555):

Yes, it does.

OpenStudy (ranga):

1 bale of wool weighs 150 kg. So how much would n bales of wool weigh?

OpenStudy (kewlgeek555):

n * 150kg

OpenStudy (ranga):

Correct. Truck weighs 4500 kg. So what is the total weight of the truck with n bales of wool?

OpenStudy (kewlgeek555):

4500?

OpenStudy (kewlgeek555):

OOPS/

OpenStudy (kewlgeek555):

4500 + 150N

OpenStudy (ranga):

Correct.

OpenStudy (ranga):

That is your function except put lowercase n. 4500 + 150n

OpenStudy (kewlgeek555):

I got confused, I am so sorry. I though it was saying that the total wieght of the truck and the bags were 4500 and ti had to equal that.

OpenStudy (kewlgeek555):

Thanks you so much! :-D

OpenStudy (ranga):

Glad to help.